OpenStudy (anonymous):

Phillip bought 12 used CDs and DVDs. CDs cost $2 each and DVDs cost $3 each. He spent $31 not including tax. How many DVDs did Phillip buy?

OpenStudy (jamesj):

Let c be the number of CDs and d the number of DVDs. Then c + d = 12 We also know that 2c + 3d = 31. Now you need to solve these two equations. For instance, from the first equation, d = 12 - c. Substitute this now into the second equation and ....

OpenStudy (anonymous):

Thank you so much.<3